OpenStudy (anonymous):

Solve for X and Y: 4X-3Y=3 5X=2Y=-25.

OpenStudy (ryan123345):

substitution or elimination?

OpenStudy (anonymous):

WHATS THAT?

OpenStudy (johnweldon1993):

you should read up on both those terms @ryan123345 said before you even approach this problem

OpenStudy (anonymous):

Idk if I ever learned that!

OpenStudy (anonymous):

eliminatioin

OpenStudy (anonymous):

**elimination

OpenStudy (anonymous):

help!(:

OpenStudy (snuggielad):

stamp thanks...she is apperantly having trouble understanding what she is learning...(no offence) this is a two step equation...i will show you wat substitution means and then i want you to give us...i mean all the people watching what you post an honest reply...that you try your best to solve it... 5(4X-3Y)+2Y=-25 try to solve that...and i will help you with the rest

OpenStudy (snuggielad):

i will also explain that

OpenStudy (snuggielad):

its addition stamp...she hit the plus and equal sign

OpenStudy (snuggielad):

key

OpenStudy (anonymous):

Ya'll are confusing me even more! )': *crying*

OpenStudy (snuggielad):

hold on stamp i will help her through this one...she only needs one persont trying to teach her at a time

OpenStudy (snuggielad):

ok...are you listening in class? because you really should be learning these words ... do you know what the mathmatic definition of distribute means? if you dont than there is really nothing i can do...stamp cant help eaither...you need to go to your teacher and ask her...there is only so much we can do without stepping over the code of conduct here...i can NOT phisically give you another answer without being banned

OpenStudy (snuggielad):

please tell me what distribute means

OpenStudy (anonymous):

it means the number outside the ( ) is multiplied by what is inside

OpenStudy (snuggielad):

YES so distribute the equation i gave above...

OpenStudy (snuggielad):

5(4X-3Y)+2Y=-25

OpenStudy (anonymous):

so it would be 20x-15y+2y=25

OpenStudy (snuggielad):

the 25 would still be negitive but yes...exactly...now solve THAT equation

OpenStudy (anonymous):

do you subtract 2y from 15y?

OpenStudy (snuggielad):

yes...

OpenStudy (anonymous):

and you get 18?

OpenStudy (anonymous):

@SnuggieLad ?

OpenStudy (snuggielad):

dont tell anyone i did this but x=-3 and y=-5 GO TO YOUR TEACHER TMR...ask her...i cant explain it in text

OpenStudy (anonymous):

m,y teacheer wont help... and this is due first thing tomorrow ):

OpenStudy (anonymous):

@SnuggieLad wuld yhu be able to explain over the phone? like call me?

OpenStudy (anonymous):

@snuggielad ?

OpenStudy (anonymous):

Is the second equation 5x - 2y = -25 or is it 5x + 2y = - 25 ?

OpenStudy (anonymous):

+

OpenStudy (anonymous):

4x - 3y = 3 --> (5)4x - 3y = 3 5x + 2y = - 25 -->(-4)5x + 2y = - 25 ----------------- 20x - 15y = 15 (result of multiplying by 5) -20x - 8y = 100 (result of multiplying by -4) ---------------- 0 - 23y = 115 -23y = 115 y = 115/-23 y = -5 now sub 5 in for y in either of the original equations 4x - 3y = 3 4x - 3(-5) = 3 4x + 15 = 3 4x = 3 - 15 4x = -12 x = -12/4 x = - 3 now lets check... 4x - 3y = 3 4(-3) - 3(-5) = 3 -12 + 15 = 3 3 = 3 (correct) answer is x = -3 and y = -5 This is the elimination method. What you have to do in this method is try to cancel out one of your variables. It can be either one, x or y. I chose to cancel out x's. To do this, you have to make your variables opposite...one has to be negative and one has to be positive. I multiplied the first equation by 5 and the 2nd equation by -4 and by doing this, I made the x's opposite so they will cancel out. This allowed me to solve for y. Once y was solved I substituted the value I got for y into the original equations and solved for x. I then plugged in my answers into the equation and checked my answers.